The business manager for the L & L Manufacturing Company makes two investments totaling $152,000. One investment pays 15% and the other investment pays 21%. If the total six-month interest income is $13,350, how much is invested at each rate?
Thanks for your help!!!

Answer by rfer(10425)

.075x+.105(152000-x)=13350
.075x+15960-.105x=13350
-.03x=-2610
x=$87,000 at 15%
152000-87000=$65,000 at 21%
remember to divide the interest in half in your calculations because investment is for six months